Exercise 3-72 (LO3-1) The American Diabetes Association recommends a blood glucose reading of less than 130 for those with Type 2 diabetes. Blood glucose measures the amount of sugar in the blood. Below are the readings for February for a person recently diagnosed with Type 2 diabetes.

102 126 113 97 112 126 109 105 124 110 105 96 106 117 126 106 123 109 114 99 114 109 120 109 109 99 113 104

a. What is the arithmetic mean glucose reading? (Round your answer to 2 decimal places.) Arithmetic mean

b. What is the median glucose reading? (Round your final answer to whole number.) Median

c. What is the modal glucose reading? Modal
Answer #1

(a)

Arithmetic mean () is given by:

(b)
Arranging numbers in ascending order, we get:

96, 97, 99, 99, 99, 102, 104, 105, 105, 106, 106, 109, 109, 109 109, 109, 110, 112, 113, 113, 114, 114, 114, 117, 120, 123, 124, 126, 126, 126

n = 30

Median = (30 + 1)/2th item = average of 15th & 16th item = (109 + 109)/2 = 109

So,

median = 109

(c)

Mode = 109

(because 109 occurs maximum number of times)